# Counterfactual Credit Guided Bayesian Optimization

### v1

@@ Abstract
Keywords Bayesian optimization, Gaussian process, Credit Assignment, Counterfactual, Sublinear regret bound
@@ 1 Introduction
Bayesian Optimization (BO) is a powerful framework for optimizing expensive-to-evaluate black-box functions, demonstrating impressive efficiency across machine learning for hyperparameter tuning and experimental design tasks [Frazier, 2018, Wang et al., 2023]. Central to BO is the balance between exploration and exploitation [Garnett, 2023]. Traditionally, BO algorithms achieve this balance implicitly via acquisition functions [Archetti and Candelieri, 2019, Wilson et al., 2018] that guide sampling decisions based on the posterior uncertainty and predicted outcomes from Gaussian Process (GP) models [MacKay et al., 1998].
However, in BO the standard exploration–exploitation trade‑off has limitations: it often wastes evaluations in suboptimal regions. BO is fundamentally a resource allocation problem. When budgets are extremely tight, finding the promising region early is more valuable than making marginal improvements later, relying solely on exploration and exploitation can be inefficient. Meanwhile, a limitation of conventional BO methods is the implicit assumption that all historical observations contribute equally to the optimization progress. In real-world landscapes, some specific samples provide more informative evidence of the global optimum than others [Koh and Liang, 2017], yet this heterogeneity is ignored and we may fail to identify key information. As a result, budget allocation is often inefficiently wasted in suboptimal areas, resulting in the budget not being reasonably allocated to areas that are more likely to have optimal values and slowing down convergence.
@@
Some existing methods add regional constraints to avoid less informative or unsafe samples and focus on promising areas [Sui et al., 2015, Acerbi and Ji, 2017, Eriksson et al., 2019, Kim et al., 2020]. However, these methods often rely on manual thresholds and do not have adaptive capabilities, and using local optimization instead of global optimization may lead to suboptimal results. In addition, some work uses expert knowledge to “Bring Human Back Into Loop” [Hvarfner et al., 2022b, 2024b], but this often requires good priors. Without accurate priors, performance can be worse than standard BO, and in real applications, we cannot have a good prior for all settings. However, these methods essentially rely on either external thresholds or expert priors. Therefore, they are unable to explicitly characterize which region is more promising from the observations itself.
To explicitly address these limitations, we propose Counterfactual Credit Guided Bayesian Optimization (CCGBO), which introduces an explicit mechanism to quantify the contribution of each historical observation using counterfactual reasoning [Harutyunyan et al., 2019, Mesnard et al., 2021, Meulemans et al., 2023]. Because credits are derived directly from the GP posterior, CCGBO does not require external priors. Specifically, counterfactual credit works as a proxy of contribution, assessing how significantly the prediction of the current optimum would degrade if a particular observation were absent, directly answering the question: “How significantly would our prediction of the current optimum deteriorate if a particular observation were absent?”. By explicitly quantifying these contributions, CCGBO moves beyond traditional exploration-exploitation trade-off to a richer three-dimensional trade-off: exploration-exploitation-importance. Integrating counterfactual credit into the BO framework empowers BO to allocate sampling resources more efficiently and target areas that truly matter, enabling precise identification and prioritization of regions critically influencing rapid convergence to the optimum. Figure 1 visually illustrates how Credit-Weighted UCB effectively directs exploration toward areas of high contribution, demonstrating more targeted queries than standard UCB [Srinivas et al., 2010]. In summary, our contributions are threefold:
@@ 2 Related Works
In recent years, there has been growing interest in leveraging data from previous experiments to improve the efficiency of BO. (1) OutlierBO. Martinez-Cantin et al. [2018] addresses the issue that spurious observations can corrupt the GP surrogate and mislead acquisition. They propose a two‐stage method using a Student-t likelihood in the GP to identify and remove outliers. (2) Continuous optimization in non-stationary environments. When the objective drifts over time, stale data may lose relevance. Bogunovic et al. [2016] extend GP-UCB to time-varying settings by periodically restarting the model to discard old data or gradually forgetting past observations via exponentially decaying weights. Building on this, Deng et al. [2022] introduces an explicit weight sequence $\omega_{t}$ into both the GP inference and acquisition computation, making recent points carry greater influence. (3) Historical data utilization in multi-task and transfer learning. In contexts where data from related tasks are available, Swersky et al. [2013] propose a multi‐task GP that learns a task covariance kernel, thereby implicitly weighting historical data according to inter‐task similarity. Feurer et al. [2018] fit a separate GP to each previous optimization run and combine their predictions with weights based on rank similarity between the source and target tasks. Hakhamaneshi et al. [2021] develop a multi-task BO, which trains a warm GP on large offline datasets and a cold GP on online samples, and dynamically trades off their suggestions based on task relevance. More recently, Mahboubi et al. [2025] introduced a point‐by‐point transfer strategy that evaluates the value of each candidate historical sample via a Gaussian mixture model and incorporates only the observations most aligned with the target task. Overall, these methods rely on pre-specified mechanisms. We instead compute per–sample weights using a counterfactual method in BO.
@@
Embedding informative priors over the location of a function’s global optimum within BO has attracted growing interest. (1) Expert‐Driven Priors. A number of works have sought to inject user beliefs directly into the acquisition process. Ramachandran et al. [2020] warp the input space according to a prior distribution over promising regions. Souza et al. [2021] and Hvarfner et al. [2022b] generalize this idea, defining a prior over the optimum that modulates any standard acquisition function. Huang et al. [2022] complement these by actively querying experts for pairwise preferences and encoding the resulting belief model into the GP prior. Wu et al. [2017] exploit the available gradient information as a form of expert knowledge, leading to the derivative‐inclusive knowledge gradient algorithm that can drastically speed up convergence. (2) Physics‐Informed and Simulator‐Based Priors. In domains where there are partial physical models, integrating these laws into BO can produce substantial gains. Khatamsaz et al. [2023] embed known material‐science relationships directly into GP kernels to optimize NiTi alloy processing, while Kobayashi et al. [2025] infuse epitaxial growth laws into the prior and demonstrate accurate extrapolation for III–V semiconductor deposition. Boltz et al. [2025] further show that using fast accelerator‐simulation models as a GP mean function permits online tuning of particle accelerators with fewer expensive measurements. (3) Functional‐Structure Priors. Beyond beliefs and data, exploiting the known structural properties of the objective can also guide BO. Li et al. [2017] inject monotonicity constraints via virtual derivative observations to enforce that the surrogate respects known increasing or decreasing trends. Kandasamy et al. [2015] address high‐dimensional spaces by assuming an additive decomposition of the objective. However, these methods all depend heavily on having correctly specified priors, whether from experts, simulators, or structural assumptions, and can perform poorly or lose robustness when such prior information is unavailable, biased, or misspecified.
@@ 3 Preliminaries
In this section, we summarize BO setup and notation, and introduce a auxiliary prior augmented acquisition.
3.1 Bayesian Optimization (BO)
@@
Problem Setting. We consider the problem of maximizing an expensive black-box objective function $f:\mathcal{X}\to\mathbb{R},~\mathbf{x}^{*}=\arg\max_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x}),$ where $\mathcal{X}\subset\mathbb{R}^{d}$ is a compact domain and each evaluation of $f(\mathbf{x})$ incurs high cost. At iteration $t$, we have collected observations $\mathcal{D}_{t}=\bigl\{(\mathbf{x}_{i},y_{i})\bigr\}_{i=1}^{t},~y_{i}=f(\mathbf{x}_{i})+\varepsilon_{i},~\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{n}^{2}),$ where $\sigma_{n}^{2}$ denotes the variance of observation noise. BO maintains a GP surrogate that, conditioned on $\mathcal{D}_{t}$, yields a posterior mean $\mu_{t}(\mathbf{x})$ and standard deviation $\sigma_{t}(\mathbf{x})$. An acquisition function $\alpha(\mathbf{x})$, built from $\mu_{t}$ and $\sigma_{t}$, then prescribes the next query point: $\mathbf{x}_{t+1}=\arg\max_{\mathbf{x}\in\mathcal{X}\setminus\{\mathbf{x}_{1},\dots,\mathbf{x}_{t}\}}\,\alpha(\mathbf{x}),$ after which we observe $y_{t+1}=f(\mathbf{x}_{t+1})+\varepsilon_{t+1},$ and update $\mathcal{D}_{t+1}=\mathcal{D}_{t}\cup\{(\mathbf{x}_{t+1},y_{t+1})\}.$ This process repeats until a stopping criterion is met, and we return $\mathbf{x}$$^{*}$ as our estimated optimizer.
@@
Acquisition Function. Given the posterior $\mathcal{N}\bigl(\mu_{t}(\mathbf{x}),\sigma_{t}^{2}(\mathbf{x})\bigr)$, BO selects the next point by maximizing an acquisition function. For the Upper Confidence Bound (UCB), we use $\mathrm{UCB}_{t}(\mathbf{x})=\mu_{t}(\mathbf{x})+\beta_{t}\,\sigma_{t}(\mathbf{x}),~\beta_{t}>0,$ and choose $\mathbf{x}_{t+1}=\arg\max_{\mathbf{x}\in\mathcal{X}\setminus\{\mathbf{x}_{1},\dots,\mathbf{x}_{t}\}}\mathrm{UCB}_{t}(\mathbf{x}).$ The new observation $(\mathbf{x}_{t+1},y_{t+1})$ is added to $\mathcal{D}_{t+1}$, and the procedure repeats until termination. Upon completion, the best observed point $\mathbf{x}$$^{*}$ is returned.
@@ 4.1 Counterfactual Credit for BO
Ideally, one would like to perform counterfactual updates for all candidate samples in a single trajectory: Even if a particular point $\mathbf{x}^{\prime}$ was not chosen at iteration $t$, we would still like to estimate “what would have happened” had we observed $f(\mathbf{x}^{\prime})$. Existing BO methods do not provide this capability, since they only update the posterior at the actually sampled points. Therefore, we propose a counterfactual credit computation as a proxy of contribution that (i) leverages the GP posterior to estimate, for each candidate $\mathbf{x}$, the hypothetical contribution to finding the optimum, and (ii) can be evaluated in closed form and with low computational overhead at each iteration. Here we remodel the calculation method of contribution from “how does choosing an observation $\mathbf{x}$ in the past affect the finding of optimal” to “given the optimal, how relevant was the choice of observation $\mathbf{x}$ in the past to achieve it?”. If a sample contributes a lot about the optimal location, it usually means that it is located in or close to the area in space where high values are most likely to appear, and it will have a larger credit accordingly. Although observations with poor results will also contribute to finding the optimal value, their contribution is smaller than that of observations near the optimal value, which will be reflected in the credit. Incorporating counterfactual credit into BO enhances the traditional optimization strategy by assigning credit to previously evaluated points, effectively guiding the optimization toward regions with higher potential.
@@ 4.2 Computation of Counterfactual Credit
where $\phi(\cdot;m,v)$ is the Gaussian density with mean $m$, variance $v$, and $\varepsilon_{c}>0$ is a small constant used to prevent division by zero and stabilize the denominator. Next, we introduce the unconditional baseline $\bar{\ell}\;=\;\frac{1}{t}\sum_{j=1}^{t}\ell_{j}.$ And define the raw counterfactual score $s_{i}\;=\;\frac{\ell_{i}}{\,\bar{\ell}+\varepsilon_{c}\,}\;-\;1,$ so that $s_{i}>0$ indicates $\mathbf{x}_{i}$ positively contributes to achieving the global proxy $Z_{t}$. Finally, we normalize and bound these scores into credits $c_{i}\in[r_{\min},\,r_{\max}]$. Let $t$ be the current number of observations. We first calculate a normalized rank $r_{i}\;=\;\frac{\bigl|\{\,j:1\leq j\leq t,\;s_{j}\leq s_{i}\}\bigr|-1}{\,t-1\,}\;\in\;[0,1].$ Then we linearly map this rank into the credit interval: $c_{i}\;=\;r_{\min}\;+\;(r_{\max}-r_{\min})\,r_{i},$ where $r_{\min}=0.1$ and $r_{\max}=1$. We set $r_{\min}=0.1$ to avoid collapsing any weight to zero, which would prematurely exclude regions. The resulting $\{c_{i}\}$ form our counterfactual credits, which can be seamlessly incorporated into augmented acquisition functions, thus realizing a three-way exploration-exploitation-credit strategy in BO. Next, we will explain how to extend the discrete credits $c_{i}$ to arbitrary candidate points.
@@ 4.3 Counterfactual Credit-Weighted Acquisition Function
11:  Obtain Counterfactual Credit-Weighted UCB $\alpha(\mathbf{x})=\bigl[(1-\lambda)+\lambda\,w_{t}(\mathbf{x})\bigr]\,\bigl[\mu(\mathbf{x})+\beta_{t}\,\sigma(\mathbf{x})\bigr]$.
12:  $\mathbf{x}_{\rm new}\leftarrow\arg\max_{\mathbf{x}\in X_{\rm cand}}\alpha(\mathbf{x})$, $y_{\rm new}\leftarrow f(\mathbf{x}_{\rm new})$.
13:  $\mathcal{D}_{t+1}\leftarrow\mathcal{D}_{t}\cup\{(x_{\mathrm{new}},\,y_{\mathrm{new}})\}$.
14:end for
15:return $\mathbf{x}^{*}=\arg\max_{(x,y)\in\mathcal{D}_{t}}y$.
@@
(1) Propagating Credits to Continuous Candidates. Given a set of candidate points $X_{\rm cand}=\{\mathbf{x}\}$, we estimate a candidate credit $c(\mathbf{x})$ by averaging the credits of its $H$ nearest neighbors in the evaluated set: $c(\mathbf{x})=\frac{1}{H}\sum_{j\in\mathcal{N}_{H}(\mathbf{x})}c_{j},$ where $\mathcal{N}_{H}(\mathbf{x})$ are the indices of the $H$ nearest $\mathbf{x}_{j}$ to $\mathbf{x}$ under Euclidean distance. This simple KNN‐based propagation yields a smooth credit field $\pi(\mathbf{x})=\tfrac{c(\mathbf{x})}{\max_{j}c_{j}}$ over $\mathcal{X}$.
(2) Credit-Weighted UCB. We then modify the UCB acquisition by incorporating the propagated credit weight:
$\mathrm{UCB}_{\rm credit}(\mathbf{x})=\bigl[(1-\lambda)+\lambda\,w_{t}(\mathbf{x})\bigr]\,\bigl[\mu(\mathbf{x})+\beta_{t}\,\sigma(\mathbf{x})\bigr],$ (4)

### v2

@@ 1 Introduction
Bayesian Optimization (BO) is a powerful framework for optimizing expensive-to-evaluate black-box functions, demonstrating impressive efficiency across machine learning for hyperparameter tuning and experimental design tasks (Frazier, 2018; Wang et al., 2023). Central to BO is the balance between exploration and exploitation (Garnett, 2023). Traditionally, BO algorithms achieve this balance implicitly via acquisition functions (Archetti and Candelieri, 2019; Wilson et al., 2018; Wang et al., 2025) that guide sampling decisions based on the posterior uncertainty and predicted outcomes from Gaussian Process (GP) models (MacKay and others, 1998).
However, in BO the standard exploration–exploitation trade‑off has limitations: it often wastes evaluations in suboptimal regions. BO is fundamentally a resource allocation problem. When budgets are extremely tight, finding the promising region early is more valuable than making marginal improvements later, relying solely on exploration and exploitation can be inefficient. Meanwhile, a limitation of conventional BO methods is the implicit assumption that all historical observations contribute equally to the optimization progress. In real-world landscapes, some specific samples provide more informative evidence of the global optimum than others (Koh and Liang, 2017), yet this heterogeneity is ignored and we may fail to identify key information. As a result, budget allocation is often inefficiently wasted in suboptimal areas, resulting in the budget not being reasonably allocated to areas that are more likely to have optimal values and slowing down convergence.
@@
Some existing methods add regional constraints to avoid less informative or unsafe samples and focus on promising areas (Sui et al., 2015; Acerbi and Ji, 2017; Eriksson et al., 2019; Kim et al., 2020). However, these methods often rely on manual thresholds and do not have adaptive capabilities, and using local optimization instead of global optimization may lead to suboptimal results. In addition, some work uses expert knowledge to “Bring Human Back Into Loop” (Hvarfner et al., 2022b, 2024b), but this often requires good priors. Without accurate priors, performance can be worse than standard BO, and in real applications, we cannot have a good prior for all settings. However, these methods essentially rely on either external thresholds or expert priors. Therefore, they are unable to explicitly characterize which region is more promising from the observations itself.
To explicitly address these limitations, we propose Counterfactual Credit Guided Bayesian Optimization (CCGBO), which introduces an explicit mechanism to quantify the contribution of each historical observation using counterfactual reasoning (Harutyunyan et al., 2019; Mesnard et al., 2021; Meulemans et al., 2023). Because credits are derived directly from the GP posterior, CCGBO does not require external priors. Specifically, counterfactual credit works as a proxy of contribution, assessing how significantly the prediction of the current optimum would degrade if a particular observation were absent, directly answering the question: “How significantly would our prediction of the current optimum deteriorate if a particular observation were absent?” By explicitly quantifying these contributions, CCGBO moves beyond traditional exploration-exploitation trade-off to a richer three-dimensional trade-off: exploration-exploitation-importance. Integrating counterfactual credit into the BO framework empowers BO to allocate sampling resources more efficiently and target areas that truly matter, enabling precise identification and prioritization of regions critically influencing rapid convergence to the optimum. Figure 1 visually illustrates how Credit-Weighted UCB effectively directs exploration toward areas of high contribution, demonstrating more targeted queries than standard UCB (Srinivas et al., 2010). In summary, our contributions are threefold:
@@ 2 Related Works
In recent years, there has been growing interest in leveraging data from previous experiments to improve the efficiency of BO. (1) OutlierBO. Martinez-Cantin et al. (2018) addresses the issue that spurious observations can corrupt the GP surrogate and mislead acquisition. They propose a two‐stage method using a Student-t likelihood in the GP to identify and remove outliers. (2) Continuous optimization in non-stationary environments. When the objective drifts over time, stale data may lose relevance. Bogunovic et al. (2016) extend GP-UCB to time-varying settings by periodically restarting the model to discard old data or gradually forgetting past observations via exponentially decaying weights. Building on this, Deng et al. (2022) introduces an explicit weight sequence $\omega_{t}$ into both the GP inference and acquisition computation, making recent points carry greater influence. (3) Historical data utilization in multi-task and transfer learning. In contexts where data from related tasks are available, Swersky et al. (2013) propose a multi‐task GP that learns a task covariance kernel, thereby implicitly weighting historical data according to inter‐task similarity. Feurer et al. (2018) fit a separate GP to each previous optimization run and combine their predictions with weights based on rank similarity between the source and target tasks. Hakhamaneshi et al. (2021) develop a multi-task BO, which trains a warm GP on large offline datasets and a cold GP on online samples, and dynamically trades off their suggestions based on task relevance. More recently, Mahboubi et al. (2025) introduced a point‐by‐point transfer strategy that evaluates the value of each candidate historical sample via a Gaussian mixture model and incorporates only the observations most aligned with the target task. Overall, these methods rely on pre-specified mechanisms. We instead compute per–sample weights using a counterfactual method.
@@
Embedding informative priors over the location of a function’s global optimum within BO has attracted growing interest. (1) Expert‐Driven Priors. A number of works have sought to inject user beliefs directly into the acquisition process. Ramachandran et al. (2020) warp the input space according to a prior distribution over promising regions. Souza et al. (2021) and Hvarfner et al. (2022b) generalize this idea, defining a prior over the optimum that modulates any standard acquisition function. Huang et al. (2022) complement these by actively querying experts for pairwise preferences and encoding the resulting belief model into the GP prior. Wu et al. (2017) exploit the available gradient information as a form of expert knowledge, leading to the derivative‐inclusive knowledge gradient algorithm that can drastically speed up convergence. (2) Physics‐Informed and Simulator‐Based Priors. In domains where there are partial physical models, integrating these laws into BO can produce substantial gains. Khatamsaz et al. (2023) embed known material‐science relationships directly into GP kernels to optimize NiTi alloy processing, while Kobayashi et al. (2025) infuse epitaxial growth laws into the prior and demonstrate accurate extrapolation for III–V semiconductor deposition. Boltz et al. (2025) further show that using fast accelerator‐simulation models as a GP mean function permits online tuning of particle accelerators with fewer expensive measurements. (3) Functional‐Structure Priors. Beyond beliefs and data, exploiting the known structural properties of the objective can also guide BO. Li et al. (2017) inject monotonicity constraints via virtual derivative observations to enforce that the surrogate respects known increasing or decreasing trends. Kandasamy et al. (2015) address high‐dimensional spaces by assuming an additive decomposition of the objective. However, these methods all depend heavily on having correctly specified priors, whether from experts, simulators, or structural assumptions, and can perform poorly or lose robustness when such prior information is unavailable, biased, or misspecified.
@@ 3 Preliminaries
In this section, we summarize BO setup and notation, and introduce an auxiliary prior augmented acquisition.
3.1 Bayesian Optimization
@@
Problem Setting. We consider the problem of maximizing an expensive black-box objective function $f:\mathcal{X}\to\mathbb{R},~\mathbf{x}^{*}=\arg\max_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x}),$ where $\mathcal{X}\subset\mathbb{R}^{d}$ is a compact domain and each evaluation of $f(\mathbf{x})$ incurs high cost. At iteration $t$, we have collected observations $\mathcal{D}_{t}=\bigl\{(\mathbf{x}_{i},y_{i})\bigr\}_{i=1}^{t},~y_{i}=f(\mathbf{x}_{i})+\varepsilon_{i},~\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{n}^{2}),$ where $\sigma_{n}^{2}$ denotes the variance of observation noise. BO maintains a GP surrogate that, conditioned on $\mathcal{D}_{t}$, yields a posterior mean $\mu_{t}(\mathbf{x})$ and standard deviation $\sigma_{t}(\mathbf{x})$. An acquisition function $\alpha(\mathbf{x})$, built from $\mu_{t}$ and $\sigma_{t}$, then prescribes the next query point: $\mathbf{x}_{t+1}=\arg\max_{\mathbf{x}\in\mathcal{X}}\,\alpha(\mathbf{x}),$ after which we observe $y_{t+1}=f(\mathbf{x}_{t+1})+\varepsilon_{t+1},$ and update $\mathcal{D}_{t+1}=\mathcal{D}_{t}\cup\{(\mathbf{x}_{t+1},y_{t+1})\}.$ This process repeats until a stopping criterion is met, and we return $\mathbf{x}$$^{*}$ as our estimated optimizer.
@@
Acquisition Function. Given the posterior $\mathcal{N}\bigl(\mu_{t}(\mathbf{x}),\sigma_{t}^{2}(\mathbf{x})\bigr)$, BO selects the next point by maximizing an acquisition function. For the Upper Confidence Bound (UCB), we use $\mathrm{UCB}_{t}(\mathbf{x})=\mu_{t}(\mathbf{x})+\beta_{t}\,\sigma_{t}(\mathbf{x}),~\beta_{t}>0,$ and choose $\mathbf{x}_{t+1}=\arg\max_{\mathbf{x}\in\mathcal{X}}\mathrm{UCB}_{t}(\mathbf{x}).$ The new observation $(\mathbf{x}_{t+1},y_{t+1})$ is added to $\mathcal{D}_{t+1}$, and the procedure repeats until termination. Upon completion, the best observed point $\mathbf{x}$$^{*}$ is returned.
@@ 4.1 Counterfactual Credit for BO
Ideally, one would like to perform counterfactual updates for all candidate samples in a single trajectory: Even if a particular point $\mathbf{x}^{\prime}$ was not chosen at iteration $t$, we would still like to estimate “what would have happened” had we observed $f(\mathbf{x}^{\prime})$. Existing BO methods do not provide this capability, since they only update the posterior at the actually sampled points. Therefore, we propose a counterfactual credit computation as a proxy of contribution that (i) leverages the GP posterior to estimate, for each candidate $\mathbf{x}$, the hypothetical contribution to finding the optimum, and (ii) can be evaluated in closed form and with low computational overhead at each iteration. Here we remodel the calculation method of contribution from “how does choosing an observation $\mathbf{x}$ in the past affect the finding of optimal” to “given the optimal, how relevant was the choice of observation $\mathbf{x}$ in the past to achieve it?”. If a sample contributes a lot about the optimal location, it usually means that it is located in or close to the area in space where high values are most likely to appear, and it will have a larger credit accordingly. Although observations with poor results will also contribute to finding the optimal value, their contribution is smaller than that of observations near the optimal value, which will be reflected in the credit. By incorporating counterfactual credit into the acquisition function, we move beyond the exploration-exploitation trade-off to a richer framework that uses the heterogeneous informativeness of samples. This enables BO to allocate sampling resources more efficiently, concentrating effort on regions that are most relevant to locating the global optimum.
@@ 4.2 Computation of Counterfactual Credit
where $\phi(\cdot;m,v)$ is the Gaussian density with mean $m$, variance $v$, and $\varepsilon_{c}>0$ is a small constant used to prevent division by zero and stabilize the denominator. Next, we introduce the unconditional baseline $\bar{\ell}\;=\;\frac{1}{t}\sum_{j=1}^{t}\ell_{j}.$ And define the raw counterfactual score $s_{i}\;=\;\frac{\ell_{i}}{\,\bar{\ell}+\varepsilon_{c}\,}\;-\;1,$ so that $s_{i}>0$ indicates $\mathbf{x}_{i}$ positively contributes to achieving the global proxy $Z_{t}$. Finally, we normalize and bound these scores into credits $c_{i}\in[r_{\min},\,r_{\max}]$. Let $t$ be the current number of observations. We first calculate a normalized rank $r_{i}\;=\;\frac{\bigl|\{\,j:1\leq j\leq t,\;s_{j}\leq s_{i}\}\bigr|-1}{\,t-1\,}\;\in\;[0,1].$ Then we linearly map this rank into the credit interval: $c_{i}\;=\;r_{\min}\;+\;(r_{\max}-r_{\min})\,r_{i},$ where $r_{\min}=0.1$ and $r_{\max}=1$. We set $r_{\min}=0.1$ to avoid collapsing any weight to zero, which would prematurely exclude regions. The resulting $\{c_{i}\}$ form our counterfactual credits, which can be seamlessly incorporated into augmented acquisition functions, thus realizing a three-way exploration-exploitation-importance strategy in BO. Next, we will explain how to extend the discrete credits $c_{i}$ to arbitrary candidate points.
@@ 4.3 Counterfactual Credit-Weighted Acquisition Function
11:  Compute UCB $\alpha(\mathbf{x})=\mu(\mathbf{x})+\beta_{t}\,\sigma(\mathbf{x})$; shift: $\tilde{\alpha}(\mathbf{x})=\alpha(\mathbf{x})-\min_{\mathbf{x}^{\prime}}\alpha(\mathbf{x}^{\prime})$.
12:  Obtain Credit-Weighted UCB $\alpha^{\mathrm{ccg}}(\mathbf{x})=\bigl[(1-\lambda)+\lambda\,w_{t}(\mathbf{x})\bigr]\,\tilde{\alpha}(\mathbf{x})$.
13:  $\mathbf{x}_{\rm new}\leftarrow\arg\max_{\mathbf{x}\in X_{\rm cand}}\alpha^{\mathrm{ccg}}(\mathbf{x})$, $y_{\rm new}\leftarrow f(\mathbf{x}_{\rm new})$.
14:  $\mathcal{D}_{t+1}\leftarrow\mathcal{D}_{t}\cup\{(\mathbf{x}_{\mathrm{new}},\,y_{\mathrm{new}})\}$.
15:end for
16:return $\mathbf{x}^{*}=\arg\max_{(\mathbf{x},y)\in\mathcal{D}_{t}}y$.
@@
(1) Propagating Credits to Continuous Candidates. Given $X_{\rm cand}$, we estimate a candidate credit $c(\mathbf{x})$ by averaging the credits of its $H$ nearest neighbors in the evaluated set: $c(\mathbf{x})=\frac{1}{H}\sum_{j\in\mathcal{N}_{H}(\mathbf{x})}c_{j},$ where $\mathcal{N}_{H}(\mathbf{x})$ are the indices of the $H$ nearest $\mathbf{x}_{j}$ to $\mathbf{x}$ under Euclidean distance. This simple KNN‐based propagation yields a smooth credit field $\pi(\mathbf{x})=\tfrac{c(\mathbf{x})}{\max_{j}c_{j}}$ over $\mathcal{X}$.
(2) Credit-Weighted UCB. To ensure that the acquisition values remain non-negative before applying credit weights, we first shift the UCB by its minimum over candidates. Let $\alpha(\mathbf{x})=\mu(\mathbf{x})+\beta_{t}\,\sigma(\mathbf{x})$ denote the standard UCB. We define the shifted acquisition as $\tilde{\alpha}(\mathbf{x})=\alpha(\mathbf{x})-\min_{\mathbf{x}^{\prime}\in X_{\mathrm{cand}}}\alpha(\mathbf{x}^{\prime}),$ so that $\tilde{\alpha}(\mathbf{x})\geq 0$ for all candidates. Since the shift is an $\mathbf{x}$-independent constant, it does not change the argmax of the unweighted acquisition. We then define the credit-weighted acquisition function:
$\mathrm{UCB}_{\rm credit}(\mathbf{x})=\bigl[(1-\lambda)+\lambda\,w_{t}(\mathbf{x})\bigr]\,\tilde{\alpha}(\mathbf{x}),$ (4)
@@
On the Role of Acquisition Shift.
A notable property of credit-weighted acquisition is that, unlike standard GP-UCB which is invariant to additive shifts, our acquisition function is sensitive to the shift value because multiplying by non-uniform weights $w_{t}(\mathbf{x})$ breaks this invariance. Subtracting the minimum UCB ensures that the lowest-UCB candidate has acquisition value zero, while higher-UCB candidates retain positive surplus values. This design amplifies the discriminative effect of credit weights: candidates in high-credit regions with large UCB surplus are strongly favored, whereas low-UCB candidates are suppressed regardless of their credit values. Without this shift, negative UCB values multiplied by small weights could yield higher acquisition values than positive UCB values multiplied by larger weights, leading to counterintuitive sampling decisions.